please help me solve the following equation 7x^(2)+4x-3=0

Answer by algebrahouse.com(1659)

7x² + 4x - 3 = 0
(7x - 3)(x + 1) = 0 {factored into two binomials}
7x - 3 = 0 or x + 1 = 0 {set each factor equal to 0}
7x = 3 or x = -1 {added 3 and subtracted 1, respectively}
x = 3/7 or x = -1 {divided left equation by 7}

For more help from me, visit: www.algebrahouse.com